OpenStudy (anonymous):

(x^4)^4 x^8/x^5 3x^2*4x^3

OpenStudy (anonymous):

(x^4)^4 =x^16

OpenStudy (anonymous):

x^8/x^5 =x^3

OpenStudy (anonymous):

3x^2*4x^3=12X^5